OpenStudy (anonymous):

The following lines are parallel. 2x + 2y = 4 and y = -x + 5 True False

OpenStudy (anonymous):

two parallel lines will have the same slope

OpenStudy (anonymous):

7x+2y=-4 2y=-7x-4 y=-7x/2 -2 slope = -7/2 PARALLEL LINE slope will be the same = -7/2 point ( 6,-3) find b y = mx +b -3 = -7/2 * 6 +b -3=-21+b b=18 the equation will be y = -7x/2 + 18 ... PERPENDICULAR slope = 2/7 point = (6,-3) Find b y=mx + b -3 = 2/7 * 6 +b -3-12/7 =b (-21-12)/7 =b -33/7 = b y = 2x/7-33/7

OpenStudy (mrnood):

you need to re-arrange the first equation into the form y = mx + b if the 2 equations then have the same value for m then they are parallel